# PACS Collaboration

### v1

@@ Appendix A Nucleon ground-state contribution in the nucleon three-point function
where the ellipsis denotes excited state contributions. The explicit formula of the ground-state contribution can be expressed by
@@
×$\times\frac{Z_{N}^{\ast}(\bm{p}^{\prime})}{2E_{N}(\bm{p}^{\prime})}e^{-E_{N}(\bm{p}^{\prime})(t_{\mathrm{sep}}-t)},$ (44)
where $Z_{N}(\bm{p})$ and $Z_{N}^{\ast}(\bm{p}^{\prime})$ correspond to the spectral weight between the initial and final states of the nucleon, and the nucleon interpolating operator, $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is defined as
@@
with ${\mathcal{O}_{\alpha}}(q)$ contains the respective nucleon form factors. For the axial-vector ($A_{\alpha}$) and pseudoscalar ($P$) currents, ${\mathcal{O}_{\alpha}}(q)$ have the following relativistically covariant decomposition in terms of the nucleon form factors as
@@
After the trace operation is performed in the right hand side of Eq. (45), the general expression of $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ for both the axial-vector and pseudoscalar cases are given by
@@
×$\times\widetilde{F}_{A}(q^{2})\left.-q_{3}(p-p^{\prime})_{\alpha}\widetilde{F}_{P}(q^{2})\right],$ (60)
@@
In the standard ratio method described in Sec. II, the quantity of $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is extracted from the ratio of the three-point function and two-point functions defined in Eq. (9). In this case, with $p_{4}=iE_{N}(\bm{q})$ and $p_{4}^{\prime}=iM_{N}$, $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is expressed as follows Sasaki:2007gw ,
@@
×$\times\left(\delta_{3\alpha}+\frac{iq_{3}}{M_{N}+E_{N}(\bm{q})}\delta_{4\alpha}\right)\widetilde{F}_{A}(q^{2})-4M_{N}q_{3}q_{\alpha}\widetilde{F}_{P}(q^{2}),$ (63)
@@ Appendix B Spectral decomposition on the nucleon three-point function
Following an argument given by Meyer-Ottnad-Schulz Meyer:2018twz , let us consider the spectral decomposition on the nucleon ground-state contribution of the nucleon three-point function in the continuum as below,
@@
×$\times\frac{Z_{N}^{\ast}(\bm{p}^{\prime})}{{p^{\prime}}^{2}+M_{N}^{2}}e^{ip_{4}^{\prime}(t_{\mathrm{sep}}-t)}.$ (66)
@@
For the $F_{A}(q^{2})$, $s_{0}$ represents a starting point of branch cut associated with a three-pion threshold as $s_{0}=9M_{\pi}^{2}$, while for the $F_{P}(q^{2})$, $s_{0}$ represents a singularity point due to the pion pole position as $s_{0}=M_{\pi}^{2}$. In the case of the $F_{P}(q^{2})$, the presence of the pion pole on the real axis in the negative $q^{2}$ region lead to
@@
which can generate the additional contribution from the expression of (64). The additional contribution for the pseudoscalar case of $\widetilde{\Lambda}^{NN}_{P}(\bm{p}^{\prime},\bm{p})$ can be also generated since the $G_{P}(q^{2})$ has the pion pole on the real axis in the negative $q^{2}$ region as well as the $F_{P}(q^{2})$.
@@
which is classified into three sections, where section (II) is responsible for the pion pole structure.
In considering the two pole contributions arising from section (II), it is worth transforming the equation as follows,
@@
where the relative signs of the two contributions differ in the case of the temporal and spatial components 444 In Ref. Meyer:2018twz , the original discussion addressed only the temporal component of the axial-vector current and omitted the spatial one. . By the standard way of evaluating residues, the integral $I_{\alpha}$ given by Eq. (69) can yield three contributions:
@@
where $q=(i(E_{N}(\bm{q})-M_{N}),\bm{q})$ and $r^{\pm}=(iE_{\pi}(\bm{q}),\pm\bm{q})$. The contribution of $I_{\alpha}^{\mathrm{A}}$ provides the usual ground-state contribution, while the other two contributions $I_{\alpha}^{\mathrm{B}}$ and $I_{\alpha}^{\mathrm{C}}$ are nothing but the pole enhanced $\pi N$ contributions, which are represented in terms of $\Delta_{+}(t,t_{\mathrm{sep}};{\bm{q}})$ and $\Delta_{-}(t,t_{\mathrm{sep}};{\bm{q}})$ in Eq. (21) and Eq. (23), respectively.
Furthermore, we observe that these coefficients explicitly contain the pion pole contributions as a function of $q^{2}$, indicating that their derivation is based on the pion pole structure of $F_{P}(q^{2})$. It is important to recall that $q^{2}=2M_{N}(E_{N}(\bm{q})-M_{N})$ under the kinematics selected for the nucleon ground-state contribution. Consequently, the coefficient of Eq. (75) contains the pion-pole contribution as
@@
Thus, in the heavy nucleon mass limit, the coefficients of Eq. (75) and Eq. (76) become the following expressions
$I_{\alpha}^{\mathrm{B}}(t,t_{\mathrm{sep}})$ $M_{N}\rightarrow\infty]{}-\frac{1}{4E_{N}(\bm{q})}\frac{r_{\alpha}^{+}}{q^{2}+M_{\pi}^{2}}e^{-(M_{N}+E_{\pi}(\bm{q}))t}e^{-M_{N}(t_{\mathrm{sep}}-t)},$ (80)
$I_{\alpha}^{\mathrm{C}}(t,t_{\mathrm{sep}})$ $M_{N}\rightarrow\infty]{}\frac{1}{4E_{N}(\bm{q})}\frac{r_{\alpha}^{-}}{q^{2}+M_{\pi}^{2}}e^{-(M_{N}+E_{\pi}(\bm{q}))t}e^{-E_{N}(\bm{q})t}e^{-(E_{N}(\bm{q})+E_{\pi}(\bm{q}))(t_{\mathrm{sep}}-t)},$ (81)
@@
Inserting Eqs. (21) and (23) into the r.h.s. of Eq. (88), we get
@@
It is worth noting that the elimination of the leading $\pi N$ contribution from ${\cal R}_{P}^{5z}(t;\bm{q})$ using Eq. (29) ensures that the generalized GT relation holds for the ground-state contribution among three form factors, namely, $F_{A}(q^{2})$, $F_{P}(q^{2})$, and $G_{P}(q^{2})$ as long as Eq. (28) is satisfied.
@@ Appendix D Simultaneous fit approach
In Sec. II.2, we proposed the leading $\pi N$ subtraction method to determine $\widetilde{F}_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$ as defined in Eq. (26) and Eq. (30), where the leading $\pi N$ contribution in the three-point functions of the axial-vector and pseudoscalar currents are subtracted by adding the specific linear combination of the time derivatives of the $A_{4}$ and $A_{i}$ correlators. Here, instead, we employ a simultaneous fit method to multiple data sets with different values of $t_{\mathrm{sep}}$ for the residual time dependence of the correlator ratios, including the leading $\pi N$ contribution, given by the terms of $\Delta_{\pm}(t,t_{\mathrm{sep}};\bm{q})$, as defined in Eq. (24).
@@
×$\times\left[\Delta E_{N}\frac{\partial_{4}\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})}{q_{i}q_{3}}+\frac{\partial_{4}{\cal R}^{5z}_{A_{4}}(t,{\bm{q}})}{iq_{3}}\right],$ (96)
@@
The resulting fitted curve and the data points (marked as filled symbols) used to fit are shown in Fig. 13 for the lowest momentum transfer (Q1) as typical examples. The fit results successfully reproduce the slight convex shapes associated with the excited-state contamination and their $t_{\mathrm{sep}}$ dependence that appeared in $W_{A_{i}}(t)$ and $W_{P}(t)$, as well as the peculiar behavior that depends linearly on the current insertion time $t$ in $W_{A_{4}}(t)$. Similarly, the corresponding $\pi N$ contribution extracted in $W_{\Delta_{+}}(t)$ is also well described by the fit results.
In Fig. 14, we next show two amplitudes of $B(\bm{q})$ and $C(\bm{q})$ appearing in the form of $\Delta_{\pm}(t,t_{\mathrm{sep}};\bm{q})$ are evaluated in four different ways. All four results are consistent with each other in both amplitudes $B(\bm{q})$ (left) and $C(\bm{q})$ (right) as a function of the four-momentum squared $q^{2}$, which is given by $2M_{N}(E_{N}(\bm{q})-M_{N})$. More interestingly, their $q^{2}$ dependence is very similar, and in both cases they grow rapidly when $q^{2}$ is small, as in the $F_{P}$ and $G_{P}$ form factors. In Fig. 15, we plot the values of $(q^{2}+M_{\pi}^{2})B(\bm{q})$ and $(q^{2}+M_{\pi}^{2})C(\bm{q})$ as a function of $q^{2}$. It has been demonstrated that the data exhibit a flat tendency in the region of $q^{2}$ that has been utilized in this study. This observation is consistent with the analytical results obtained in Appendix B, which show that the corresponding $B(\bm{q})$ and $C(\bm{q})$ possess the pion-pole nature as a function of $q^{2}$.
Intuitively, if the nucleon form factor has a pion pole structure such as $F_{P}(q^{2})$ and $G_{P}(q^{2})$, the contribution from the excited state should be relatively suppressed in comparison with the ground-state contribution in the low $q^{2}$ region. This is simply because the magnitude of the form factor increases rapidly with lower $q^{2}$. In fact, however, the $\pi N$-state contamination is much more pronounced in both cases of $F_{P}(q^{2})$ and $G_{P}(q^{2})$. The underlying reason for this reversal trend is now straightforward: the amplitudes of the terms contributed by the leading $\pi N$ excited states are also enhanced by the pion-pole nature, in conjunction with the ground state contributions, $F_{P}(q^{2})$ and $G_{P}(q^{2})$. In other words, the ground state contribution and the leading $\pi N$ contribution are both enhanced by the pion pole effect, causing their magnitudes to compete with each other at lower $q^{2}$.
@@
Figure 13: Typical results of the simultaneous fits on $W_{A_{i}}(t)$ (upper-left), $W_{P}(t)$ (upper-right), $W_{A_{4}}(t)$ (lower-left) and $W_{\Delta_{+}}(t)$ (lower-right) at the lowest momentum transfer for the $160^{4}$ lattice ensemble. In each panel, there are three data sets obtained with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles). The red curves and the gray bands represent the resulting fitted curves, while the data points used to fit are marked as filled symbols.
@@ Appendix E Test for pion-pole structure
First of all, the axial Ward-Takahashi identity for the nucleon three-point function, where both the ground state and excited state contributions are entangled, is verified in Ref. Tsuji:2023llh . This observation suggests that the ground-state contribution and the excited-state contribution must separately satisfy the axial Ward-Takahashi identity. For the ground state, this is nothing but the generalized GT relation. Therefore, once the leading $\pi N$ contribution on the axial-vector three-point function is identified as $\Delta_{\pm}(t)$ term, the leading $\pi N$ contribution on the pseudoscalar three-point function can be described by $\Delta_{P}(t)=Z_{A}B_{0}\Delta_{+}(t)$, where the factor $Z_{A}B_{0}$ is determined to hold the axial Ward-Takahashi identity for the leading $\pi N$ contribution. Using the terms of $\Delta_{+}(t)$ and $\Delta_{P}(t)$, the true ground-state contribution, $\widetilde{F}_{P}(q^{2})$ ($\widetilde{G}_{P}(q^{2})$) and $\widetilde{F}^{\mathrm{std}}_{P}(q^{2})$ ($\widetilde{G}^{\mathrm{std}}_{P}(q^{2})$) evaluated by the standard ratio method are related as below

### v2

@@ Appendix A Nucleon ground-state contribution in the nucleon three-point function
where the ellipsis denotes excited-state contributions. The explicit formula of the ground-state contribution can be expressed by
@@
$\times\frac{Z_{N}^{\ast}(\bm{p}^{\prime})}{2E_{N}(\bm{p}^{\prime})}e^{-E_{N}(\bm{p}^{\prime})(t_{\mathrm{sep}}-t)},$ (44)
where $Z_{N}(\bm{p})$ ($Z_{N}^{\ast}(\bm{p}^{\prime})$) corresponds to the spectral weight between the nucleon interpolating operator and the initial (final) state of the nucleon, $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is defined as
@@
with ${\mathcal{O}_{\alpha}}(q)$ containing the respective nucleon form factors. For the axial-vector ($A_{\alpha}$) and pseudoscalar ($P$) currents, ${\mathcal{O}_{\alpha}}(q)$ have the following relativistically covariant decomposition in terms of the nucleon form factors as
@@
After the trace operation is performed in the right-hand side of Eq. (45), the general expression of $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ for both the axial-vector and pseudoscalar cases are given by
@@
$\times\widetilde{F}_{A}(q^{2})\left.-q_{3}(p-p^{\prime})_{\alpha}\widetilde{F}_{P}(q^{2})\right],$ (60)
@@
In the standard ratio method described in Sec. II, the quantity of $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is extracted from the ratio of the three-point function and two-point functions defined in Eq. (9). In our choice of the kinematics with $p_{4}=iE_{N}(\bm{q})$ and $p_{4}^{\prime}=iM_{N}$, $\widetilde{\Lambda}^{NN}_{J}(\bm{p}^{\prime},\bm{p})$ is expressed as follows Sasaki:2007gw ,
@@
$\times\left(\delta_{3\alpha}+\frac{iq_{3}}{M_{N}+E_{N}(\bm{q})}\delta_{4\alpha}\right)\widetilde{F}_{A}(q^{2})-4M_{N}q_{3}q_{\alpha}\widetilde{F}_{P}(q^{2}),$ (63)
@@ Appendix B Spectral decomposition on the nucleon three-point function
Following an argument given by Meyer-Ottnad-Schulz Meyer:2018twz , let us consider the spectral decomposition on the nucleon ground-state contribution of the nucleon three-point function in the continuum as follows:
@@
$\times\frac{Z_{N}^{\ast}(\bm{p}^{\prime})}{{p^{\prime}}^{2}+M_{N}^{2}}e^{ip_{4}^{\prime}(t_{\mathrm{sep}}-t)}.$ (66)
@@
For the $F_{A}(q^{2})$, $s_{0}$ represents a starting point of branch cut associated with a three-pion threshold as $s_{0}=9M_{\pi}^{2}$, while for the $F_{P}(q^{2})$, $s_{0}$ represents a singularity point due to the pion-pole position as $s_{0}=M_{\pi}^{2}$. In the case of the $F_{P}(q^{2})$, the presence of the pion pole on the real axis in the negative $q^{2}$ region leads to
@@
which can generate the additional contribution from the expression of (64). The additional contribution for the pseudoscalar case of $\widetilde{\Lambda}^{NN}_{P}(\bm{p}^{\prime},\bm{p})$ also can be generated, since the $G_{P}(q^{2})$ has the pion pole on the real axis in the negative $q^{2}$ region as well as the $F_{P}(q^{2})$.
@@
which is classified into three sections, where section (II) is responsible for the pion-pole structure.
In considering the two pole contributions arising from section (II), it is worth transforming the equation as follows:
@@
where the relative signs of the two contributions differ in the case of the temporal and spatial components.  [^5] In Ref. Meyer:2018twz , the original discussion addressed only the temporal component of the axial-vector current and omitted the spatial one. By the standard way of evaluating residues, the integral $I_{\alpha}$ given by Eq. (69) can yield three contributions:
@@
where $q=(i(E_{N}(\bm{q})-M_{N}),\bm{q})$ and $r^{\pm}=(iE_{\pi}(\bm{q}),\pm\bm{q})$. The contribution of $I_{\alpha}^{\mathrm{A}}$ provides the usual ground-state contribution, while the other two contributions $I_{\alpha}^{\mathrm{B}}$ and $I_{\alpha}^{\mathrm{C}}$ are nothing but the pole-enhanced $\pi N$ contributions, which are represented in terms of $\Delta_{+}(t,t_{\mathrm{sep}};{\bm{q}})$ and $\Delta_{-}(t,t_{\mathrm{sep}};{\bm{q}})$ in Eqs. (21) and (23), respectively.
Furthermore, we observe that these coefficients explicitly contain the pion-pole contributions as a function of $q^{2}$, indicating that their derivation is based on the pion-pole structure of $F_{P}(q^{2})$. It is important to recall that $q^{2}=2M_{N}(E_{N}(\bm{q})-M_{N})$ under the kinematics selected for the nucleon ground-state contribution. Consequently, the coefficient of Eq. (75) contains the pion-pole contribution as
@@
Thus, in the heavy nucleon mass limit, the coefficients of Eqs. (75) and (76) become the following expressions
$I_{\alpha}^{\mathrm{B}}(t,t_{\mathrm{sep}})$ $\xrightarrow[M_{N}\rightarrow\infty]{}-\frac{1}{4E_{N}(\bm{q})}\frac{r_{\alpha}^{+}}{q^{2}+M_{\pi}^{2}}e^{-(M_{N}+E_{\pi}(\bm{q}))t}e^{-M_{N}(t_{\mathrm{sep}}-t)},$ (80)
$I_{\alpha}^{\mathrm{C}}(t,t_{\mathrm{sep}})$ $\xrightarrow[M_{N}\rightarrow\infty]{}\frac{1}{4E_{N}(\bm{q})}\frac{r_{\alpha}^{-}}{q^{2}+M_{\pi}^{2}}e^{-(M_{N}+E_{\pi}(\bm{q}))t}e^{-E_{N}(\bm{q})t}e^{-(E_{N}(\bm{q})+E_{\pi}(\bm{q}))(t_{\mathrm{sep}}-t)},$ (81)
@@
Inserting Eqs. (21) and (23) into the right-hand side of Eq. (88), we get
@@
It is worth noting that the elimination of the leading $\pi N$ contribution from ${\cal R}_{P}^{5z}(t;\bm{q})$ using Eq. (29) ensures that the generalized GT relation holds for the ground-state contribution among three form factors, namely, $F_{A}(q^{2})$, $F_{P}(q^{2})$, and $G_{P}(q^{2})$, as long as Eq. (28) is satisfied.
@@ Appendix D Simultaneous fit approach
In Sec. II.2, we proposed the leading $\pi N$ subtraction method to determine $\widetilde{F}_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$ as defined in Eqs. (26) and (30), where the leading $\pi N$ contribution in the three-point functions of the axial-vector and pseudoscalar currents are subtracted by adding the specific linear combination of the time derivatives of the $A_{4}$ and $A_{i}$ correlators. Here, instead, we employ a simultaneous fit method to multiple datasets with different values of $t_{\mathrm{sep}}$ for the residual time dependence of the correlator ratios, including the leading $\pi N$ contribution, given by the terms of $\Delta_{\pm}(t,t_{\mathrm{sep}};\bm{q})$, as defined in Eq. (24).
@@
$\times\left[\Delta E_{N}\frac{\partial_{4}\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})}{q_{i}q_{3}}+\frac{\partial_{4}{\cal R}^{5z}_{A_{4}}(t,{\bm{q}})}{iq_{3}}\right],$ (96)
@@
The resulting fitted curve and the data points (marked as filled symbols) used to fit are shown in Fig. 13 for the lowest momentum transfer (Q1) as typical examples. The fit results successfully reproduce the slight convex shapes associated with the excited-state contamination and their $t_{\mathrm{sep}}$ dependence that appear in $W_{A_{i}}(t)$ and $W_{P}(t)$, as well as the peculiar behavior that depends linearly on the current insertion time $t$ in $W_{A_{4}}(t)$. Similarly, the corresponding $\pi N$ contribution extracted in $W_{\Delta_{+}}(t)$ is also well described by the fit results.
In Fig. 14, we next show two amplitudes of $B(\bm{q})$ and $C(\bm{q})$ appearing in the form of $\Delta_{\pm}(t,t_{\mathrm{sep}};\bm{q})$ are evaluated in four different ways. All four results are consistent with each other in both amplitudes $B(\bm{q})$ (left) and $C(\bm{q})$ (right) as a function of the four-momentum squared $q^{2}$, which is given by $2M_{N}(E_{N}(\bm{q})-M_{N})$. More interestingly, their $q^{2}$ dependence is very similar, and in both cases they grow rapidly when $q^{2}$ is small, as in the $F_{P}$ and $G_{P}$ form factors. In Fig. 15, we plot the values of $(q^{2}+M_{\pi}^{2})B(\bm{q})$ and $(q^{2}+M_{\pi}^{2})C(\bm{q})$ as a function of $q^{2}$. It has been demonstrated that the data points exhibit a flat tendency in the region of $q^{2}$ that has been utilized in this study. This observation is consistent with the analytical results obtained in Appendix B, which show that the corresponding $B(\bm{q})$ and $C(\bm{q})$ possess the pion-pole nature as a function of $q^{2}$.
Intuitively, if the nucleon form factor has a pion-pole structure such as $F_{P}(q^{2})$ and $G_{P}(q^{2})$, the contribution from the excited state should be relatively suppressed in comparison with the ground-state contribution in the low $q^{2}$ region. This is simply because the magnitude of the form factor increases rapidly with lower $q^{2}$. In fact, however, the $\pi N$-state contamination is much more pronounced in both cases of $F_{P}(q^{2})$ and $G_{P}(q^{2})$. The underlying reason for this reversal trend is now straightforward: the amplitudes of the terms contributed by the leading $\pi N$ excited states are also enhanced by the pion-pole nature, in conjunction with the ground-state contributions, $F_{P}(q^{2})$ and $G_{P}(q^{2})$. In other words, the ground-state contribution and the leading $\pi N$ contribution are both enhanced by the pion-pole effect, causing their magnitudes to compete with each other at lower $q^{2}$.
@@
Figure 13: Typical results of the simultaneous fits on $W_{A_{i}}(t)$ (upper-left), $W_{P}(t)$ (upper-right), $W_{A_{4}}(t)$ (lower-left) and $W_{\Delta_{+}}(t)$ (lower-right) at the lowest momentum transfer for the $160^{4}$ lattice ensemble. Each panel contains three datasets obtained with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and 1919 (circles). The red curves and the gray bands represent the resulting fitted curves, while the data points used to fit are marked as filled symbols.
@@ Appendix E Test for pion-pole structure
First of all, the axial Ward-Takahashi identity for the nucleon three-point function, where both the ground-state and excited-state contributions are entangled, is verified in Ref. Tsuji:2023llh . This observation suggests that the ground-state contribution and the excited-state contribution must separately satisfy the axial Ward-Takahashi identity. For the ground state, this is nothing but the generalized GT relation. Therefore, once the leading $\pi N$ contribution on the axial-vector three-point function is identified as $\Delta_{\pm}(t)$ term, the leading $\pi N$ contribution on the pseudoscalar three-point function can be described by $\Delta_{P}(t)=Z_{A}B_{0}\Delta_{+}(t)$, where the factor $Z_{A}B_{0}$ is determined to hold the axial Ward-Takahashi identity for the leading $\pi N$ contribution. Using the terms of $\Delta_{+}(t)$ and $\Delta_{P}(t)$, the true ground-state contribution, $\widetilde{F}_{P}(q^{2})$ ($\widetilde{G}_{P}(q^{2})$) and $\widetilde{F}^{\mathrm{std}}_{P}(q^{2})$ ($\widetilde{G}^{\mathrm{std}}_{P}(q^{2})$) evaluated by the standard ratio method are related as follows: